A typical firm in​ long-run equilibrium in an industry with identical firms has a cost function given by

​C(q)=4,900+2q2.

What is the equilibrium​ price?


1
Expert's answer
2021-10-13T10:04:58-0400

In long run equilibrium, Price = Average cost (AC) = Marginal cost (MC)


"AC =\frac {C(q)} {Q} = (\frac {4,900} {Q} ) + 4q"


"MC =\frac {dC(q)} {dq} = 4q"


Equating AC & MC,


"(\frac {4,900} {Q} ) + 2q = 4q"


"2q = \frac {4,900} {q}"


"2q2 = 4,900"


"q2 = 2450"


"q=49.5"


Price = MC = 4q = 4 x 49.5 = 198